OpenStudy (anonymous):

175 as a fraction in lowest terms

OpenStudy (anonymous):

175?

OpenStudy (anonymous):

i mean 176/100

OpenStudy (lalaly):

88/50 = 44/25

jimthompson5910 (jim_thompson5910):

Divide each component by the GCF 4 to get \[\large \frac{176}{100}=\frac{44}{25}\]

OpenStudy (anonymous):

176/100 = = 88/50 = 44/25

OpenStudy (anonymous):

which?

OpenStudy (lalaly):

44/25

OpenStudy (anonymous):

it's 44/25 =]